OpenStudy (anonymous):

Essay question: show all work. Find the difference: (-5x^2+2x-5)-(7x^2-2x-3)

OpenStudy (anonymous):

for the second half of the question you need to distribute the negative sign. so it becomes: (-7x^2+2x+3) so the full equation now is (-5x^2+2x-5)-(-7x^2+2x-3) Then you subtract life terms. So it is: -5x^2-(-7x^2)=2x^2 then: 2x+2x=4x then: -5-(-3)=2 so it would come out to: 2x^2+4x-2

OpenStudy (anonymous):

ok so what about the first half

OpenStudy (anonymous):

of the equation?

OpenStudy (anonymous):

yes does it stay the same

OpenStudy (anonymous):

yes. the negative sign only affects the second half of the equation.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so it thats the whole problem or is more needed to be done

OpenStudy (anonymous):

BUT you only distribute the negative sign when its a subtraction problem, not addition. and yes. the answer at the bottom is the full answer.

OpenStudy (anonymous):

ok so if it is addition you leave it as it is

OpenStudy (anonymous):

yesss. i did the equation horizontally, but it might be easier to do it vertically. look at the problem like this\[-5x^2+2x-5\over7x^2-2x-3\] think of the line in the middle as the line that tells you to subtract. since all the like terms are lined up it makes it easier.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

make more sense?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

Now i have a question about another problem

OpenStudy (anonymous):

ok shoot